# The dimension of sparse random graph orders

### v1

@@ Abstract
A random graph order is a partial order obtained from a random graph on $[n]$ by taking the transitive closure of the adjacency relation. The dimension of the random graph orders from random bipartite graphs $\mathcal{B}(n,n,p)$ and from $\mathcal{G}(n,p)$ were previously studied when $p=\Omega(\log n/n)$, and there is a conjectured phase transition in the sparse range $p=O(1/n)$. We investigate the dimensions of the partial orders arising from $\mathcal{B}(n,n,p)$ and $\mathcal{G}(n,p)$ in this sparse range. We also prove an upper bound on the dimension of general partial orders based on decompositions of partial orders into suborders, which are of independent interest.
@@ 1 Introduction
a.a.s. $\dim P_{\mathcal{B}(n,n,p)}>(\delta pn\log pn)/(1+\delta p\log pn)$.
They also asked many questions regarding $\dim P_{\mathcal{B}(n,n,p)}$, such as whether $\dim P_{\mathcal{B}(n,n,p)}$ is unimodal as $p$ increases? When does $\dim P_{\mathcal{B}(n,n,p)}$ become at least $k$ for constant $k$? What happens to $\dim P_{\mathcal{B}(n,n,p)}$ when $p$ is very close to 1? These questions remain open. As comparison, Bollobás and Brightwell [5] proved the following regarding $\dim P_{\mathcal{G}(n,p)}$, improving earlier bounds given by Albert and Frieze [1], and extending the range of $p$ under study .
@@
In view of Theorem 1, there is a phase transition around $p=1/\log n$, and $\dim P_{\mathcal{B}(n,n,p)}\sim n$ for $1/\log n\ll p<1-n^{-1+\epsilon}$. There must be at least another phase transition at some $p>1-n^{-1+\epsilon}$ since $\dim P_{\mathcal{B}(n,n,p)}$ eventually drops back to two. How this happens has not been understood well yet. Theorem 2 shows that the behaviour of $\dim P_{\mathcal{G}(n,p)}$ is quite different from $\dim P_{\mathcal{B}(n,n,p)}$: there are phase transitions somewhere around $p=\log n/n$ and somewhere around $p=1/\log n$. Bollobás and Brightwell predicted “a third transition phase, as $p$ decreases through $c/n$, since the underlying random graph changes from being almost all in one component, to being a collection of small trees”. It is known that $\dim P_{\mathcal{G}(n,p=c/n)}\leq 4$ when $c<1$ (see Theorems 3 and 4 below), since $\mathcal{G}(n,p=c/n)$ is a collection of small trees and unicyclic components. When $c>1$, $\mathcal{G}(n,p=c/n)$ contains a giant component whose order is linear in $n$, and has many cycles and other complicated structures. For instance, for every fixed $k\geq 1$, $\mathcal{G}(n,p=c/n)$ has a subgraph with minimum degree at least $k$, provided that $c$ is sufficiently large (depending on $k$). Does “a third transition phase” appears in the way that right after $c$ passes 1, $\dim P_{\mathcal{G}(n,p)}$ immediately blows up to $\omega(1)$? Indeed, Bollobás and Brightwell specifically asked in [5] what happens to $\dim P_{\mathcal{G}(n,p)}$ if $p=c/n$ with fixed $c>1$.
@@ 1.1 Preliminary and notations
Let $P$ be a poset on a finite set of elements $V$. Without loss of generality we denote $V$ by $[n]=\{1,2,\ldots,n\}$ where $n$ is the number of elements in $V$. We always use $<_{P}$ to denote the order in $P$, and use $<$ to denote the order in natural numbers. For $x,y\in V$, we use $x||y$ to denote that $x$ and $y$ are incomparable in $P$.
A realiser of a partial order $P$ on a finite set of elements $V$ is a set of total orders =$\mathcal{R}=\{L_{1},\ldots,L_{k}\}$ on $V$ satisfying the following conditions.
@@
How about poset whose cover graph has a unique cycle? Bollobás and Brightwell [5] conjectured that such posets have dimension at most three. The best known upper bound is four, which follows from the upper bound on the dimension of a poset whose cover graph is outerplanar [10]. There is also a short proof provided by Brightwell (from personal communications). We include his proof below for completeness. Given a total ordering $L$ on a set of elements $V$ and $S\subseteq V$, let $L|_{S}$ denote the total ordering on $S$ obtained by restricting $L$ on $S$. Given a subset of elements $S$ of $P$, let $U_{P}[S]$ and $D_{P}[S]$ denote the up-set and the down-set of $S$, i.e.
@@
This is a proof provided by Brightwell (via personal communications). Let $G$ be the unicycle cover graph, and let $C$ be the cycle in $G$. Let $x$ be a maximal element in $C$. Let $A$ be the set of neighours of $x$ in $G$ that are greater than $x$ in $P$, and let $u_{1},\ldots,u_{k}$ be an enumeration of the vertices in $A$. Let $W_{i}$ be the component of $G-x$ containing $u_{i}$ for $1\leq i\leq k$. Let $W[x]=\{x\}\cup(\cup_{i=1}^{k}W_{i})$, and let $R=V(G)\setminus W[x]$. It is readily seen that both the cover graphs of the subposets of $P[W[x]]$ and $P[R]$ are both acyclic, and hence each of them has dimension at most three. Moreover, $U(x)\subseteq W[x]$ and $D(x)\subseteq R$. Let $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ be a realiser for $P[W[x]]$ and $\{\tau_{1},\tau_{2},\tau_{3}\}$ be a realiser for $P[R]$. Let $\hat{L}$ be an arbitrary total ordering on the elements in $P$. Construct four linear orderigs on the set of elements in $P$ as follows.
@@
[…]
@@ 1.2 Main results
There are two well known upper bounds for the dimension of a poset $P$. The first is the width of $P$, which is the size of a largest anti-chain. The second upper bound is given by Fürhedi and Kahn [11] in terms of the maximum number of elements comparable to a given element in the set (Theorem 5). As mentioned earlier, this bound is proved to be almost optimal, as Erdös, Kierstad and Trotter [9] showed an almost-matching lower bound on the dimension of $P_{\mathcal{B}(n,n,p)}$, when $p\geq\log^{1+\epsilon}n/n$. Bollobás and Brightwell proved a similar phenomenon in $P_{\mathcal{G}(n,p)}$ for $p=\Omega(\log n/n)$: the dimension is close to the minimum of the width, and the Fürhedi-Kahn bound (Theorem 2). This is no longer the case for sparser random graph orders. Consider $P_{\mathcal{G}(n,p=c/n)}$ for an example. The minimum of the above two bounds would be around $\log n\log^{2}\log n$. However, although the maximum number of comparable elements to a given element can be as large as $\log n$, most elements are only comparable to $O(e^{c})$ other elements, and the number of “problematic” elements who are comparable to more than $ce^{Kc}$ other elements are bounded by $e^{-Kc}n$ (see Corollary 30). It is then reasonable to have a careful study of the sub-posets induced by these “problematic” elements, and study how they impact the dimension of the original poset. We prove the following theorems which upper bounds the dimension of a poset by the sum of certain sub-posets. Although these bounds are specifically tailored to help to bound the dimension of sparse random graph orders, we believe that they are of independent interest, as we are not aware of bounds of this nature in the literature.
@@
Using Theorem 7 we prove the following bounds on the dimension of $P_{\mathcal{B}(n,n,p)}$ where $p=c_{n}/n$. We assume that $c_{n}\geq 1$ and $c_{n}\leq\log^{2}n$, since the dimension of $P_{\mathcal{B}(n,n,p)}$ for $p<1/n$ is well understood by Theorem 3, and the case $p=\Omega(\log^{2}n/n)$ has already been studied by Erdös, Kierstad and Trotter [9].
@@
    a.a.s. $\dim P_{\mathcal{B}(n,n,p)}\geq(1-o(1))\frac{1}{2\alpha_{c}}$;
@@
Remark. $\alpha_{c_{n}}\sim 2\log c_{n}/c_{n}$ as $c_{n}\to\infty$ (see Lemma 16), and thus $\dim P_{\mathcal{B}(n,n,p=c/n)}\geq(1-o_{c}(1))c/4\log c$ as $c\to\infty$. On the other hand, for small values of $c_{n}$, $\alpha_{c_{n}}$ can be numerically computed. The figure below on the left is a plot of $1/2\alpha_{c}$ for $302<c<$ 30. From the plot the reader can see that the lower bounds obtained for small values of $c$ (below 20) are smaller than the trivial lower bound two. We stress that the merit of Theorem 9, together with the upper and lower bounds for $\dim P_{\mathcal{G}(n,p)}$ in Theorems 10 and 11, is about the asymptotic behaviour of $\dim P_{\mathcal{B}(n,n,p)}$ when $c$ goes large. In fact, the implicit constant coefficients in the upper bounds (in both Theorems 9 and 11) are inherited from the implicit constant in the Füredi-Kahn bound, and thus these bound do not yield useful information when $c$ is small. Obtaining nontrivial lower or upper bounds on $P_{\mathcal{B}(n,n,p)}$ for small values of $c$ remains open. See Open Problem 13 below.
[…]
@@
Let $p=c_{n}/n$ where $c_{n}\geq 1$ and $c_{n}=o(\log n)$. Then, there exist $\gamma:=\gamma(c)>0$ where $\lim_{c\to\infty}\gamma(c)=1$ such that a.a.s. $\dim P_{\mathcal{G}(n,p)}=O(c_{n}^{4}e^{c_{n}})$.
@@
Open problem 13.
What is the largest positive integer $I$ that satisfies the following property: for all fixed $c>1$, a.a.s. $\dim P_{\mathcal{G}_{n}}\geq I$, where $\mathcal{G}_{n}\sim\mathcal{G}(n,p=c/n)$ or $\mathcal{G}_{n}\sim\mathcal{B}(n,n,p=c/n)$?
@@ 1.3 Connections to causal set theory
Causal set theory is a theoretical framework that models spacetime as a fundamentally discrete, combinatorial structure. In contrast to general relativity, which treats spacetime as a smooth manifold, this approach—motivated by ideas from quantum gravity and quantum mechanics—posits that spacetime consists of indivisible ”atoms” of events, organized into a locally finite poset; see [17]. The partial order in this poset encodes the causal relationships between events. A central challenge in causal set theory is to generate such posets in a way that is both physically meaningful and mathematically rigorous. One of the most influential approaches to this problem is the classical sequential growth (CSG) model. Among the models studied, the random graph model $P_{\mathcal{G}(n,p)}$ serves as a primary example of a CSG model within this framework; see [7] for work connecting the studies of $P_{\mathcal{G}(n,p)}$ with developments in causal set theory.
A key concept in the study of causal sets is the causal representation of a poset in Minkowski space. Roughly speaking, the goal is to embed a poset $P$ into the $d$-dimensional Minkowski space such that the order relation corresponds to lightcone containment—that is, $x<y$ in $P$ if and only if the event $x$ lies in the causal past of $y$ in the embedding. The Minkowski dimension of $P$is the smallest dimension $d$ for which such a representation exists. We refer interested readers to [15] for more details. While both the (order-theoretic) dimension and the Minkowski dimension measure the complexity of a poset, they do so in fundamentally different ways. Studying the Minkowski dimension of the random poset $P_{\mathcal{G}(n,p)}$ remains an intriguing and open problem, linking combinatorics, geometry, and the foundations of spacetime.
2 Proof of Theorem 7
@@
Now the first assertion of Theorem 7 follows by Lemma 6. For the second assertion, consider $P^{\prime}$, the dual of $P$. Then, $U_{P^{\prime}}(S_{B})=D_{P}(S_{B})\subseteq A$ and $U_{P^{\prime}}(B\setminus S_{B})=D_{P}(B\setminus S_{B})\subseteq A$. Thus, the second assertion follows by the first assertion of Theorem 7 by symmetry.
We prove Claim 14 next.
Proof of Claim 14 Let $x,y$ be elements in $P$ with $x\nleq_{P}y$.

### v2

# The dimension of sparse and co-sparse random graph orders

@@ Abstract
A random graph order is a partial order obtained from a random graph on $[n]$ by taking the transitive closure of the adjacency relation. The dimension of the random graph orders from random bipartite graphs $\mathcal{B}(n,n,p)$ and from $\mathcal{G}(n,p)$ were previously studied when $p=\Omega(\log n/n)$ and when $p$ is not too close to 1. There is a conjectured phase transition in the sparse range at $p=1/n$. In this paper, we investigate this conjectured phase transition and estimate the dimension of the partial orders arising from $\mathcal{B}(n,n,p)$ and $\mathcal{G}(n,p)$ when $p=O(1/n)$. For the random bipartite order, we additionally estimate its dimension in the co-sparse regime, thereby closing all previously open ranges of $p$. Finally, we establish a general upper bound on the dimension of partial orders based on their decompositions into suborders, a result that is of independent interest.
@@ 1 Introduction
$a.$s. $\dim P_{\mathcal{B}(n,n,p)}>(\delta pn\log pn)/(1+\delta p\log pn)$.
Their lower bound is tight up to a logarithmic factor, by the upper bound given by Fürhedi and Kahn [12]; see Theorem 5 below. In the same paper [10], they also asked many questions regarding $\dim P_{\mathcal{B}(n,n,p)}$, such as whether $\dim P_{\mathcal{B}(n,n,p)}$ is unimodal as $p$ increases? When does $\dim P_{\mathcal{B}(n,n,p)}$ become at least $k$ for constant $k$? What happens to $\dim P_{\mathcal{B}(n,n,p)}$ when $p$ is very close to 1? These questions remained open. As a comparison, Bollobás and Brightwell [6] proved the following regarding $\dim P_{\mathcal{G}(n,p)}$, improving earlier bounds given by Albert and Frieze [2], and extending the range of $p$ under study.
@@
In view of Theorem 1, there is a phase transition around $p=1/\log n$. Afterwards, for $1/\log n\ll p<1-n^{-1+\epsilon}$, $\dim P_{\mathcal{B}(n,n,p)}\sim n$. There must be at least another phase transition at some $p>1-n^{-1+\epsilon}$ since $\dim P_{\mathcal{B}(n,n,p)}$ eventually drops back to two. In this paper, we capture this phase transition by determining the asymptotic order of the dimension of $P_{\mathcal{B}(n,n,p)}$ for all $p\geq 1-n^{-1+1/7}$, addressing a question (see [10, Open problem (5.8)]) of Erdős, Kierstead and Trotter: see Theorem 9 below. Theorem 2 shows that the behaviour of $\dim P_{\mathcal{G}(n,p)}$ is quite different from $\dim P_{\mathcal{B}(n,n,p)}$: there are phase transitions around $p=\log n/n$ and then around $p=1/\log n$. Bollobás and Brightwell predicted “a third transition phase, as $p$ decreases through $c/n$, since the underlying random graph changes from being almost all in one component, to being a collection of small trees”. It is known that $\dim P_{\mathcal{G}(n,p=c/n)}\leq 4$ when $c<1$ (see Theorems 3 and 4 below), since $\mathcal{G}(n,p=c/n)$ is a collection of small trees and unicyclic components. When $c>1$, $\mathcal{G}(n,p=c/n)$ contains a giant component whose order is linear in $n$, and has many cycles and other complicated structures. For instance, for every fixed $k\geq 1$, $\mathcal{G}(n,p=c/n)$ has a subgraph with minimum degree at least $k$, provided that $c$ is sufficiently large (depending on $k$). Does “a third transition phase” appears in the way that right after $c$ passes 1, $\dim P_{\mathcal{G}(n,p)}$ immediately blows up to $\omega(1)$? Indeed, Bollobás and Brightwell specifically asked in [6] what happens to $\dim P_{\mathcal{G}(n,p)}$ if $p=c/n$ with fixed $c>1$.
@@ 1.1 Preliminary and notations
Let $P$ be a poset on a finite set of elements $V$. Without loss of generality we denote $V$ by $n]=\{1,2,\ldots,n\}$ where $n$ is the number of elements in $V$. We always use $_{P}$ to denote the order in $P$, and use << to denote the order in natural numbers. For $x,y\in V$, we use $x||y$ to denote that $x$ and $y$ are incomparable in $P$.
A realiser of a partial order $P$ on a finite set of elements $V$ is a set of total orders $\mathcal{R}=\{L_{1},\ldots,L_{k}\}$ on $V$ satisfying the following conditions.
@@
How about poset whose cover graph has a unique cycle? Bollobás and Brightwell [6] conjectured that such posets have dimension at most three. Abram and Segovia recently reported solving this conjecture [1]. The best known upper bound before [1] is four, which follows from the upper bound on the dimension of a poset whose cover graph is outerplanar [11]. There is also a short proof provided by Brightwell (from personal communications). Since his proof is short and self-contained, we include it below for completeness. Given a total ordering $L$ on a set of elements $V$ and $S\subseteq V$, let $L|_{S}$ denote the total ordering on $S$ obtained by restricting $L$ on $S$. Given a subset of elements $S$ of $P$, let $U_{P}[S]$ and $D_{P}[S]$ denote the up-set and the down-set of $S$, i.e.
@@
This is a proof provided by Brightwell (via personal communications). Let $G$ be the unicycle cover graph, and let $C$ be the cycle in $G$. Let $x$ be a maximal element in $C$. Let $A$ be the set of neighours of $x$ in $G$ that are greater than $x$ in $P$, and let $u_{1},\ldots,u_{k}$ be an enumeration of the vertices in $A$. Let $W_{i}$ be the component of GxG-x containing $u_{i}$ for $1\leq i\leq k$. Let $W[x]=\{x\}\cup(\cup_{i=1}^{k}W_{i})$, and let $R=V(G)\setminus W[x]$. It is readily seen that both the cover graphs of the subposets of $P[W[x]]$ and $P[R]$ are acyclic, and hence each of them has dimension at most three. Moreover, $U(x)\subseteq W[x]$ and $D(x)\subseteq R$. Let $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ be a realiser for $P[W[x]]$ and $\{\tau_{1},\tau_{2},\tau_{3}\}$ be a realiser for $P[R]$. Let $\hat{L}$ be an arbitrary total ordering on the elements in $P$. Construct four linear orderigs on the set of elements in $P$ as follows.
@@
Note that the notation in the first inequality above means that first ordering elements in $R$ according to $\sigma_{i}$, followed by elements in $W[x]$ ordered according to $\tau_{i}$. It is straightforward to verify that $\{\pi_{i}:\ 1\leq i\leq 4\}$ is a realiser for $P$. ∎
@@ 1.2 Main results
There are two well known upper bounds for the dimension of a poset $P$. The first is the width of $P$, which is the size of a largest anti-chain. The second upper bound is given by Fürhedi and Kahn [12] in terms of the maximum number of elements comparable to a given element in the set (Theorem 5). As mentioned earlier, this bound is almost optimal in view of Theorem 1, since it matches, up to a logarithmic factor, the lower bound on the dimension of $P_{\mathcal{B}(n,n,p)}$, when $p\geq\log^{1+\epsilon}n/n$ (and $p$ is not too close to 1). Bollobás and Brightwell proved a similar phenomenon in $P_{\mathcal{G}(n,p)}$ for $p=\Omega(\log n/n)$: the dimension is close to the minimum of the width, and the Fürhedi-Kahn bound (Theorem 2). This is no longer the case for sparser random graph orders. Consider $P_{\mathcal{G}(n,p=c/n)}$ for an example. The minimum of the above two bounds would be around $\log n\log^{2}\log n$. However, although the maximum number of comparable elements to a given element can be as large as $\log n$, most elements are only comparable to $O(e^{c})$ other elements, and the number of “problematic” elements who are comparable to more than $ce^{Kc}$ other elements are bounded by $e^{-Kc}n$ (see Corollary 35). It is then reasonable to have a careful study of the sub-posets induced by these “problematic” elements, and study how they impact the dimension of the original poset. We prove the following theorems which upper bounds the dimension of a poset by the sum of certain sub-posets. Although these bounds are specifically tailored to help to bound the dimension of sparse random graph orders, we believe that they are of independent interest, as we are not aware of bounds of this nature in the literature.
@@
Now we turn to random graph orders. We start from addressing a question of Erdős, Kierstead and Trotter regarding the dimension of $P_{\mathcal{B}(n,n,p)}$ when $p$ is very close to 1.
Theorem 9.
Let q:=1pq:=1-p.
  1. (a)
    If $q=O(1/n)$ and $q=\omega(1/n^{2})$ then a.a.s. $\dim P_{\mathcal{B}(n,n,p)}=\Theta(n^{2}q)$, and $n-\dim P_{\mathcal{B}(n,n,p)}=\Omega(n)$.
  2. (b)
    If $q=\omega(1/n)$ and $q\leq n^{-1+1/7}$ then a.a.s. $\dim P_{\mathcal{B}(n,n,p)}=(1-o(1))n$.
Remark. The vanishing error term $o(1)$ in part (b) (cf. the $O(1/\log n)$ error in Theorem 1) can in fact be made explicit from our proof, although we did not pursue such a quantification. Our main message here is that there is a phase transition from $q=\omega(1/n)$ to $q=O(1/n)$, at which point the dimension of the bipartite partial order ceases to be asymptotic to $n$. We excluded the uninteresting case $q=O(1/n^{2})$ from part (a) since in that regime $\bar{B}$ contains $O_{p}(1)$ isolated edges and immediately $\dim P_{\mathcal{B}(n,n,p)}=O_{p}(1)$.
Next we turn to the sparse regime. Using Theorem 7 we prove the following bounds on the dimension of $P_{\mathcal{B}(n,n,p)}$ where $p=c_{n}/n$. We assume that $c_{n}\geq 1$ and $c_{n}\leq\log^{2}n$, since the dimension of $P_{\mathcal{B}(n,n,p)}$ for $p<1/n$ is well understood by Theorem 3, and the case $p=\Omega(\log^{2}n/n)$ has already been studied by Erdös, Kierstad and Trotter [10].
@@
    $a.$s. $\dim P_{\mathcal{B}(n,n,p)}\geq(1-o(1))\frac{1}{2\alpha_{c_{n}}}$;
@@
Remark. $\alpha_{c_{n}}\sim 2\log c_{n}/c_{n}$ as $c_{n}\to\infty$ (see Lemma 17), and thus $\dim P_{\mathcal{B}(n,n,p=c/n)}\geq(1-o_{c}(1))c/4\log c$ as $c\to\infty$. On the other hand, for small values of $c_{n}$, $\alpha_{c_{n}}$ can be numerically computed. The figure below on the left is a plot of $1/2\alpha_{c}$ for $2<c<30$. From the plot the reader can see that the lower bounds obtained for small values of $c$ (below 20) are smaller than the trivial lower bound two. We stress that the merit of Theorem 10, together with the upper and lower bounds for $\dim P_{\mathcal{G}(n,p)}$ in Theorems 11 and 12, is about the asymptotic behaviour of $\dim P_{\mathcal{B}(n,n,p)}$ when $c$ goes large. In fact, the implicit constant coefficients in the upper bounds (in both Theorems 10 and 12) are inherited from the implicit constant in the Füredi-Kahn bound, and thus these bound do not yield useful information when $c$ is small. Obtaining nontrivial lower or upper bounds on $P_{\mathcal{B}(n,n,p)}$ for small values of $c$ remains open. See Open Problem 14 below.
Next we consider the dimension of $P_{\mathcal{G}(n,p)}$. Let $Li_{2}(z):=-\int_{0}^{z}y^{-1}\log(1-y)dy$ denote the dilogarithm function, also known as the Spencer’s function.
@@
Let $p=c_{n}/n$ where $c_{n}\geq 1$ and $c_{n}=o(\log n)$. Then, a.a.s. $\dim P_{\mathcal{G}(n,p)}=O(c_{n}^{4}e^{c_{n}})$.
@@
Open problem 14.
What is the largest positive integer $I$ that satisfies the following property: for all fixed $c>1$, $a.$s. $\dim P_{\mathcal{G}_{n}}\geq I$, where $\mathcal{G}_{n}\sim\mathcal{G}(n,p=c/n)$ or $\mathcal{G}_{n}\sim\mathcal{B}(n,n,p=c/n)$?
@@ 1.3 Connections to causal set theory
Causal set theory is a theoretical framework that models spacetime as a fundamentally discrete, combinatorial structure. In contrast to general relativity, which treats spacetime as a smooth manifold, this approach—motivated by ideas from quantum gravity and quantum mechanics—posits that spacetime consists of indivisible “atoms” of events, organized into a locally finite poset; see [18]. The partial order in this poset encodes the causal relationships between events. A central challenge in causal set theory is to generate such posets in a way that is both physically meaningful and mathematically rigorous. One of the most influential approaches to this problem is the classical sequential growth (CSG) model. Among the models studied, the random graph model $P_{\mathcal{G}(n,p)}$ serves as a primary example of a CSG model within this framework; see [8] for work connecting the studies of $P_{\mathcal{G}(n,p)}$ with developments in causal set theory.
A key concept in the study of causal sets is the causal representation of a poset in Minkowski space. Roughly speaking, the goal is to embed a poset $P$ into the $d$-dimensional Minkowski space such that the order relation corresponds to the lightcone containment—that is, $x<y$ in $P$ if and only if the event $x$ lies in the causal past of $y$ in the embedding. The Minkowski dimension of $P$ is the smallest dimension $d$ for which such a representation exists. We refer interested readers to [16] for more details. While both the (order-theoretic) dimension and the Minkowski dimension measure the complexity of a poset, they do so in fundamentally different ways. Studying the Minkowski dimension of the random poset $P_{\mathcal{G}(n,p)}$ remains an intriguing and open problem, linking combinatorics, geometry, and the foundations of spacetime.
2 Proof of Theorem 7.
@@
Now the first assertion of Theorem 7 follows by Lemma 6 and Claim 15. For the second assertion, consider $P^{\prime}$, the dual of $P$. Then, $U_{P^{\prime}}(S_{B})=D_{P}(S_{B})\subseteq A$ and $U_{P^{\prime}}(B\setminus S_{B})=D_{P}(B\setminus S_{B})\subseteq A$. Thus, the second assertion follows from the first assertion of Theorem 7 by symmetry.
It only remains to prove Claim 15.
Proof of Claim 15. Let $x,y$ be elements in $P$ with $x\nleq_{P}y$. We consider two cases.